OpenStudy (anonymous):

A triangle has sides which are 6, 8, and 10 inches long. Which applied equation will prove that this triangle is or is not a right triangle

OpenStudy (anonymous):

use tan for all of the sides so tanA= 8/6, tanB=10/6, tanC=8/10 and since none of them are equal to 0/1 (which is tan90) that proves that none of the angles are 90 (which would mean that it is a right triangle

OpenStudy (anonymous):

let a=6, b=8, c=10. a^2+b^2=c^2. 6^2+8^2=10^2 36+64=100, therefore it is a right triangle pythagorean theorem is only valid for right triangles.